OpenStudy (anonymous):

This one is a doozy.. X/x^2-4 (greater than or equal to sign) 0. How would I solve ?

OpenStudy (lgbasallote):

\[\frac x{x^2 - 4} \ge 0\] ???

OpenStudy (anonymous):

yes that is the problem

OpenStudy (lgbasallote):

cross multiply..what do you get?

OpenStudy (anonymous):

1x=0

OpenStudy (anonymous):

?

OpenStudy (lgbasallote):

= ?

OpenStudy (lgbasallote):

where did you get =?

OpenStudy (anonymous):

1x = 0

OpenStudy (lgbasallote):

should be \(\ge\) not =

OpenStudy (lgbasallote):

your initial sign was \(\ge\) so you can't change it

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so my answer is correct ? or is there more to do ?

OpenStudy (lgbasallote):

nothing more

OpenStudy (lgbasallote):

or if you want to be meticulous like other "mathematicians" note that x cannot be 2 or -2

OpenStudy (anonymous):

Uh... @lgbasallote , that answer isn't right.

OpenStudy (lgbasallote):

you sure?

OpenStudy (anonymous):

yep:)

OpenStudy (anonymous):

Oh ok, just making sure because I am supposed to solve this as an inequality and show a sign chart as well expressing in interval notation

OpenStudy (lgbasallote):

then why don't you show how it's done, oh great one?

OpenStudy (anonymous):

It's \(x\in (-2, 0]\cup (2,\infty)\) Well, that's what I'm getting... let me check.

OpenStudy (lgbasallote):

i did say x cannot be 2 or -2

OpenStudy (anonymous):

According to Mathematica, that seems to be the case. And, I'm not trying to make you look bad or anything, jeez, it's a mistake, and it happens, don't take it personally. Check the boundaries yourself.

OpenStudy (lgbasallote):

you did read what i was saying right?

OpenStudy (anonymous):

Yes, but, it's supposed to be: \[ x>2\\ -2<x\le 0 \]

OpenStudy (anonymous):

I'd look at where x^3-4x is positive and negative...

OpenStudy (anonymous):

simplest way I think...

OpenStudy (lgbasallote):

i didn't say any signs did i? i said x cannot be 2 or -2

OpenStudy (anonymous):

So how does one get to this conclusion? If you dont mind me asking.

OpenStudy (lgbasallote):

and i was also insulting the mathematicians

OpenStudy (anonymous):

multiply both sides by (x^2-4 )^2

OpenStudy (anonymous):

I'm actually not sure what the 'standard' way is to look at these sorts of inequalities... I've forgotten maybe. Someone else might have a more kosher method.

OpenStudy (anonymous):

\[\frac x{x^2 - 4} \ge 0\] \[\frac x{(x-2)(x+2)} \ge 0\] changes sign at \(-2,0,2\) so only need to check on interval. say \(x<-2\) i pick \(x=-3\) we get \[\frac{-3}{(-3)^2-4}=\frac{-3}{5}\] which is negative so it is negative, positive, negative, positive on the respective intervals \[(-\infty,-2), (-2,0), (0,2), (2,\infty)\]

OpenStudy (anonymous):

Here's a counter example to your solution: \[ x=1\ge0\\ \frac{1}{1-4}=-\frac{1}{3}\not \ge\; 0 \]

OpenStudy (anonymous):

counter example to what?

OpenStudy (lgbasallote):

you're a mathematician @LolWolf ...you know that greater than and less than are under "inequalities" => not equal to and again i repeat...i was insulting the meticulousness of mathemeticians

OpenStudy (anonymous):

(@lgbasallote 's not yours, @satellite73 , haha)

OpenStudy (anonymous):

oh, ok

OpenStudy (lgbasallote):

oh

OpenStudy (lgbasallote):

anyway..if you don't need the domain of this then plain \(x \ge 0\) is fine

OpenStudy (lgbasallote):

if it's a computer you're trying to please

OpenStudy (anonymous):

Thank you. The other question I have goes like this : a^5/3 * b^2/3 * c^3 / C^-2 * a^2/3 * b^-4/3 .

OpenStudy (anonymous):

add/subtract the exponents on like bases eg. : a^(2/3) * a^(5/3) = a^(7/3)

OpenStudy (anonymous):

for example on (a) would leave denominator 3 alone and just add the numnerators?

OpenStudy (anonymous):

Hope you're not asking about adding fractions :P

OpenStudy (anonymous):

Nope!

OpenStudy (anonymous):

My answer is A^1 * b^6/3 * c^5

OpenStudy (anonymous):

would this be correct?